OpenStudy (anonymous):

Charlie takes classes at both Westside Community College and Pinewood Community College. At Westside, class fees are $98 per credit hour, and at Pinewood, class fees are $115 per credit hour. Charlie is taking a combined total of 13 credit hours at the two schools. Suppose that he is taking w credit hours at Westside. Write an expression for the combined total dollar amount he paid for his class fees. total paid (in dollars) =

OpenStudy (anonymous):

@sangya21 @kah.x ????? do you know

OpenStudy (anonymous):

98w+115p=total amount paid.

OpenStudy (anonymous):

it doesnt give you enough info to actually solve it

OpenStudy (anonymous):

@kah.x is right

OpenStudy (anonymous):

its says it wrng

OpenStudy (anonymous):

@sangya21 @kah.x was the 13 supposed to be used

OpenStudy (anonymous):

w+p =13

OpenStudy (anonymous):

yeah, but you cant really put that into the current equation because theres not enough information to solve it

OpenStudy (anonymous):

they wont let me put a equal sign

OpenStudy (anonymous):

it was 98w+115(13-w) @sangya21

OpenStudy (anonymous):

thats also a way to put. Sorry for the wrong

OpenStudy (anonymous):

Yesterday, Rachel went on a bike ride. Her average speed was 10 miles per hour. Today, she went on another ride, this time averaging 14 miles per hour. In the two days, she biked for a combined total time of 10 hours. Let x be the number of hours she biked yesterday. Write an expression for the combined total number of miles she biked in the two days.

OpenStudy (anonymous):

its okay can u help with that one @sangya21 @kah.x

OpenStudy (anonymous):

you're so helpful

OpenStudy (anonymous):

we are given speed and time for yesterday speed = 10mph time = x hours distance = speed * time d1= 10x d2 = distance on day 2 d2 = 14(10-x) 10x+14(10-x)

OpenStudy (anonymous):

@uri @Becki @ChasingMaggie @sangya21 The sum of the four angle measures of any convex quadrilateral is 360 degrees. Suppose that a convex quadrilateral has angle measures of 60 , +8y3 , and −6y7 degrees. Write an expression for the degree measure of the fourth angle.